Solve the equation abx=cy for each of the following
x
a
c
i know that i have to do something with both sides but ivd tried and tried and i just dont understand...please be very simple

Answer by Nate(3500)

for x:
abx=cy the equation
(abx)/(ab) = (cy)/(ab) divide each side by 'ab'
x = (cy)/(ab)
remember: (a)/(a) = 1 when a is not zero
for a:
abx=cy
(a)/(bx)=(cy)/(bx)
a = (cy)/(bx)
for c:
abx=cy
(abx)/(y) = (cy)/(y)
(abx)/(y) = c